OpenStudy (prettygirl_shynice):

How many feet are there in 35 yards (there are 3 feet in one yard)like can u explain how i get the answer

OpenStudy (prettygirl_shynice):

@jigglypuff314

jimthompson5910 (jim_thompson5910):

multiply \(\Large 35 \ \text{yards}\) by the conversion factor \(\Large \frac{3 \ \text{feet}}{1 \ \text{yard}}\)

OpenStudy (prettygirl_shynice):

im confused

OpenStudy (prettygirl_shynice):

but i dont get it but dont tell me answer

jigglypuff314 (jigglypuff314):

it's the opposite idea here, I'll try to draw it pretend these are 2 yards 

jimthompson5910 (jim_thompson5910):

Notice how the "yards" unit cancels out \[\Large 35 \ \text{yards} \times \frac{3 \ \text{feet}}{1 \ \text{yard}}\] \[\Large 35 \ \cancel{\text{yards}} \times \frac{3 \ \text{feet}}{1 \ \cancel{\text{yard}}}\]

OpenStudy (prettygirl_shynice):

yea

OpenStudy (prettygirl_shynice):

okay im listening

OpenStudy (prettygirl_shynice):

yhall there

jimthompson5910 (jim_thompson5910):

do you see how I used the conversion factor?

OpenStudy (prettygirl_shynice):

what that

jigglypuff314 (jigglypuff314):